OpenStudy (anonymous):

I have to verify that this is true using trig identities.. Let me type it real quick

OpenStudy (anonymous):

\[((\sin \theta)/(1-\cos \theta))-(\sin \theta \cos \theta)/(1+\cos \theta)=\csc \theta(1+\cos ^2\theta)\]

OpenStudy (anonymous):

take common denominator u can solve it easily

OpenStudy (mertsj):

\[\frac{(\sin \theta)(1+\cos \theta)}{(1-\cos \theta)(1+\cos \theta)}-\frac{(\sin \theta \cos \theta)(1-\cos \theta)}{(1-\cos \theta)(1+\cos \theta)}\]

OpenStudy (anonymous):

let teta be U [sin U(1+cosU)-(1-cosU)(sinUcosU)]/1-cos^2 U =[sinU+cosUsinU-sinUcosU+sinUcos^2U]/sin^2 U =csc U+cscUcos^2U =cscU(1+cos^2U)

OpenStudy (mertsj):

\[\frac{1+\cos \theta}{\sin \theta}-\frac{\cos \theta(1-\cos \theta)}{\sin \theta}=\frac{1+\cos \theta-\cos \theta+\cos ^{2}\theta}{\sin \theta}\]\[\frac{(\sin \theta)(1+\cos \theta)}{\sin ^{2}\theta}-\frac{(\sin \theta \cos \theta)(1-\cos \theta)}{\sin ^{2}\theta}\]

OpenStudy (mertsj):

\[\frac{1+\cos \theta}{\sin \theta}-\frac{(\cos \theta)(1-\cos \theta)}{\sin \theta}=\frac{1+\cos \theta-\cos \theta+\cos ^{2}\theta}{\sin \theta}\]

OpenStudy (mertsj):

\[\csc \theta(1+\cos ^{2}\theta)\]